If there were 600 people at a play, the total receipts were 9,800 pesos, and the admission price was 20 pesos for adults and 10 pesos for children, how many adults and how many children were in attendance? .....thank you
Answer by mananth(16949)

Adult x Numbers
Child ------------ y Numbers
x+y=600 .............1
20x+10y=9800 .............2
Eliminate y
multiply (1)by -10
Multiply (2) by 1
-10x-10y=-6000
20x+10y=9800
Add the two equations
10x= 3800
/10
x=380
plug value of x in (1)
x+y=600
380+y=600
y=600 -380
y=220